OpenStudy (anonymous):

Find the zeros (roots) of the quadratic equation x^2−8x+12=0 A. -4, 6 B. 4, -4 C. 2, 6 D. -2, 3

OpenStudy (anonymous):

Let us factorize the given quadratic equation: \[x^2−8x+12=0 \rightarrow x^2−6x-2x+12=0 \] \[\rightarrow x(x−6)-2(x-6)=0 \rightarrow (x−6)(x-2)=0 \] i.e. either (x-2)=0 or (x−6) =0 i.e. x=2 or x=6 Thus the zeros (roots) of the quadratic equation are 2, 6 Hence, option C. 2, 6 is the correct answer. @LaBlueEyes2013